### v1

# Evaluating Long Range Dependency Handling in Code Generation Models using Multi-Step Key Retrieval

Abstract
As language models support larger and larger context sizes, evaluating their ability to make effective use of that context becomes increasingly important. We analyze the ability of several code generation models to handle long range dependencies using a suite of multi-step key retrieval tasks in context windows up to 8k tokens in length. The tasks progressively increase in difficulty and allow more nuanced evaluation of model capabilities than tests like the popular needle-in-the-haystack test. We find that performance degrades significantly (up to 2x) when a function references another function that is defined later in the prompt. We also observe that models that use sliding window attention mechanisms have difficulty handling references further than the size of a single window. We perform simple prompt modifications using call graph information to improve multi-step retrieval performance up to 3x. Our analysis highlights different facets of long-context performance and is suggestive of prompt construction strategies for code completion tools.
1 Introduction
Long context inference is an increasingly important differentiator of LLMs, with model releases supporting larger and larger context windows [1, 2, 3, 4, 5]. This is enabled by advances in efficient attention implementation such as FlashAttention [6], Grouped-Query Attention [7] and Paged Attention [8], as well as scaling attention by introducing sparsity via windowing [9, 10].
Applications, such as in-editor code completion from tools like GitHub’s Copilot [11] and SourceGraph’s Cody [12], benefit from long context support as they leverage retrieval-augmented generation strategies [13] to incorporate code snippets from across the user’s project into a single prompt. This allows completions to be driven by the user’s context rather than just by what the model learned at training time. Works such as Zhang et al. [14] and Shrivastava et al. [15] have proposed various approaches to improve cross-repository code completion.
These applications rely heavily on the model’s ability to effectively use everything in its context window, and thus evaluation of long context usage has been a topic of interest in the community. Khandelwal et al. [16] showed that early LSTM-based language models only roughly modelled information from more distant tokens. Recently, the “needle-in-the haystack” test, which measures recall for chat-style LLMs [17] has grown popular, and a parallel key-retrieval method has been used by Rozière et al. [18] to measure recall in auto-complete style code generation models. We believe that desirable long context inference also includes the ability to reason over multiple pieces of information in the input, and needle-in-haystack evaluation approaches do not capture a model’s ability to do this. In the paper we contribute:
  1. 1.
    A set of multi-step key retrieval tasks that progressively increase in difficulty and allow evaluation of long-range dependency handling and basic reasoning through indirection.  [^1] We open source the code to generate these tasks at https://github.com/apple/ml-key-retrieval-code-tasks.
  2. 2.
    Empirical evaluation of several open source code generation models that reveal architectural choices that affect ability to handle long-range dependencies. We also discover that the order of function declarations has a large effect on model ability to complete these tasks.
  3. 3.
    Preliminary recommendations for improving multi-step recall that rely only on prompt modification using information obtainable via existing non-LLM techniques. These include rewriting code to bring dependencies closer together and adding annotations of function dependencies in the form of comments.
We focus on examining how well long-range dependencies are handled in auto-complete style code generation models. While instruction-tuned/chat-style models are an increasingly popular mode of LLM interaction, we believe that auto-complete models are still relevant for applications that need frequent, low latency generation of code.
2 Related Work
Symbolic reasoning: Zhang et al. [19] study neural networks’ ability to handle indirection by introducing the Pointer Value Retrieval (PVR) task. They train models to take in a sequence of tokens (typically digits), where the first token acts as pointer to the value to be retrieved. The synthetic nature of this task provides full control over the complexity of the problem. Abnar et al. [20] extend this work to add recursive steps to the PVR task, evaluate it in the context of vision transformers, and present a new architecture that supports adaptive compute.
Multi-hop QA: Min et al. [21] study a popular multi-hop reasoning benchmark (HotpotQA) and find that many of the questions can be answered with single hop reasoning. They find many questions embed contextual info that make the task easier, and that weak distractors may make picking the right answer ‘obvious’ without need for the desired reasoning steps. Chen and Durrett [22] further explore dataset design choices to induce multi-hop reasoning.
Long context retrieval: Liu et al. [23] Examine instruction-tuned LLM performance for synthetic key-retrieval tasks as well as multi-document QA and describe a “lost-in-the-middle” phenomenon where information becomes harder to retrieve if it is not near the beginning or end of the prompt.
Long context dependencies: Yu et al. [24] present a code completion benchmark named CoderEval, where problem solutions have varying levels of dependencies on code in the surrounding context. They find that model performance for functions with dependencies is significantly worse than performance for standalone functions. In the context of natural-language Kuratov et al. [25], Yuan et al. [26] and Levy et al. [27] develop benchmarks to evaluate how well language models extract and reason over distributed facts in the presence of noise text. Hsieh et al. [28] develop a rich set of tasks to evaluate long context performance with different dependency structures in natural language.
Chain-of-thought prompting: Wei et al. [29] and Kojima et al. [30] explore prompting strategies to induce multi-step reasoning in instruction tuned models. They find that models perform better on various tasks when prompted to output intermediate steps in the reasoning chain.
While most of the existing literature focuses on chat-style natural language generation, our work focuses on non-instruction-tuned models in the context of autocomplete-style code generation, which are not able to make use of ‘scratchpad’ methods such as chain-of-thought to improve their reasoning. We situate our work in between the PVR work of Zhang et al. [19], which is highly controllable but uses a fully synthetic task, and the code-with-dependencies benchmarking work exemplified by Yu et al. [24], which is more realistic but also makes it more difficult to run controlled experiments that allow discovery of specific failure modes.
3 Long Context Multi-Step Key Retrieval
3.1 Task Design
To study long-range dependency handling we propose four multi-step key retrieval tasks of increasing complexity (one-step, two-step, three-step, and concatenation retrieval). These extend the key-retrieval task in Rozière et al. [18] and test models’ ability to integrate multiple pieces of information spread throughout a long context window to make a completion.
(a) One Step
def key():
return "xdfgew"
assert key() ==
(b) Two Step
def value():
return "xdfgew"
def key():
return value()
assert key() ==
(c) Three Step
def value_2():
return "xdfgew"
def value_1():
return value_2()
def key():
return value_1()
assert key() ==
(d) Concatenation
def value_1():
return "xdfgew"
def value_2():
return "asdahj"
def key():
return value_1() + value_2()
assert key() ==
Figure 1: Key retrieval tasks with increasing levels of difficulty. Function names and return values are randomized in the actual prompt. See Appendix F for a complete example.
One-step retrieval is equivalent to the key-retrieval task in [18]. We differ from their design by using random strings to construct function names rather than fixed function names, and return values that are string literals rather than integer literals. Using a simple template system, we first generate a key function that returns a random string. The model is then asked to complete an assert statement on the return value of this function. All return strings that are $10$ characters long and all function names are between $13$ and $20$ characters long.
In two-step retrieval, the key function calls a value function that returns the string. Three-step retrieval adds an additional function call between the key function and the value function that returns the string. In the concatenation retrieval task the key function calls two value functions and returns the concatenation of their returned strings Figure 1.
To turn each of these into a long context problem we insert varying amounts irrelevant code into the context window, we detail this in Section 3.3. In all cases the assert statement is on the result of the key function and is placed at the end of the prompt.
3.2 Avoiding Reliance on Parametric Knowledge and Trivial Solutions
To make sure the model is not solely relying on parametric knowledge (i.e., knowledge stored in the weights) to solve the task, we construct function names from two or three random sequences of lowercase characters or digits, joined with underscores (e.g., zcxjdz_309521_xcdgfp). Return values are random strings of lowercase characters (e.g., pczjdfeyxc). This makes it extremely unlikely that such functions exist in the model’s training data, and therefore to complete the task successfully, the model must use the information from the prompt.
In our early experiments we observed that models may produce trivial solutions such as assert foo() == foo(). While these are technically correct , such solutions prevent us from testing the models’ ability to retrieve information from the context. To prevent this, we guide the decoding of tokens to make sure that the response starts as a string literal. As we sample output tokens, we use the prefix_allowed_tokens functionality in the HuggingFace transformers library [31] to ensure that the output starts with any number of spaces followed by a single or a double quote. Once an initial quote mark is produced, generation proceeds unrestricted.
3.3 Long Context Construction
To turn the tasks described above into long context inference problems we add irrelevant snippets to the context window from two sources. Motivated by findings in Min et al. [21] and Chen and Durrett [22] that show that the presence of good distractor functions are important when measuring multi-step reasoning, we first generate a number of synthetic distractor functions that use the same template as the the key and value functions that contain the task information. Then we sample standalone Python functions from the HumanEval dataset [32] to fill out the context window to our desired size. Algorithm 1 describes this dataset generation process in more detail.
Algorithm 1 Generate Long Context Retrieval Tasks
$n_{k}\leftarrow$ number of unique key functions
$n_{d}\leftarrow$ number of synthetic distractor functions
$n_{t}\leftarrow$ maximum number of tokens
$n_{p}\leftarrow$ maximum number of position combinations
repeat $n_{k}$ times
[…]
end
This generates a dataset that allows us to compare the effect of the following variables: position of task related snippets within the input context, relative order of the task-related snippets, and the spread of task related snippets. We vary these factors while keeping the irrelevant functions used constant for each key function we generate. To aid reproducibility we fix the random seed during generation so that the prompt set can easily be re-generated.
In our experiments, we set $n_{k}$ to $100$ for one-step task, and $20$ for other tasks; $n_{d}$ to 00, $1$, and $5$ for three distractor conditions; $n_{p}$ to $\infty$, $150$, $50$, $50$ for one-step, two-step, three-step, and concatenation tasks respectively (this results in a maximum of 6000 prompts for each condition). We set $n_{t}$ to 2k, 4k and 8k tokens.
3.4 Models
We evaluate these tasks on five open source models: StarCoderBase-1B, StarCoderBase-7B, StarCoderBase-15.5B [33], StarCoder2-7B [34], and Mistral-7B-v0.1 [4]. The StarCoder family of models are competitive on code generation benchmarks and available in a number of different parameter counts. Mistral is a general purpose LLM that is nonetheless competitive on code generation benchmarks. We use implementations from the HuggingFace transformers library [31].
Table 1: Models evaluated.
Model HumanEval [^2] Mistral-7B was trained with 8192 context size but reports support for up to 131k tokens. Max Context Size33footnotemark: 3 Sliding Window Size
Mistral-7B 30.5 8192 / 131072 4096
StarCoder2-7B 35.4 16384 4096
StarCoderBase-1B 15.1 8192 N/A
StarCoderBase-7B 30.5 8192 N/A
StarCoderBase-15.5B 29.3 8192 N/A
11footnotetext: HumanEval scores for StarCoder* models are from Lozhkov et al. [34], StarcoderBase-1B scores are from https://huggingface.co/bigcode/starcoderbase-1b. Mistral-7B scores are from Jiang et al. [4].
We are primarily interested in seeing how the performance of a given model changes as task difficulty (i.e., number of hops needed) and context size varies. We evaluate multiple models to see how consistent effects are across multiple models.
3.5 Results
3.5.1 Overall Task Performance
Our primary metric is accuracy@k, which is similar to the pass@k metric introduced in Chen et al. [32], the only difference is instead of running unit tests, we simply check if the string literal produced is the expected string. We compute accuracy@3 over 10 generations for each input prompt. Hyperparameters for generation are in Appendix C
(a) Task performance by model.
(b) Task performance by prompt length.
Figure 2: Overall task performance. For detailed scores, see Table 4 in appendix.
Figure 3: Effect of number of distractors by task variant. For detailed scores, see Table 5 in appendix.
Task difficulty: 2(a) shows that tasks increase in difficulty in the following order: one-step, two-step, three-step, and concatenation. We note particularly weak performance on the concatenation task. Another factor affecting difficulty is the number of distractor functions. For the more difficult tasks we see a with a smooth degradation in performance as distractors are added (Figure 3). We focus the rest of our analysis on the five distractor condition as it allows us to see a good variation in task performance.
Context size: In 2(b) we observe a small performance drop as context sizes increases for StarCoderBase-1B, StarCoderBase-7B, and StarCoderBase-15. For StarCoder2-7B and Mistral-8B, which both use sliding window attention, a large drop in performance occurs when moving from 4k to 8k context length, we hypothesise that this is related to the size of the sliding window used and will explore this in more detail in Section 3.5.2.
Incorrect responses: We did some manual inspection of the incorrect generations from models and note a number of distinct failure modes. Firstly a noticeable amount of incorrect generations are return values of distractor functions. Secondly we see completely incorrect answers, such as "$3$", or "text". The third failure mode we notice is partially correct results, particularly for the concatenation task where models often fail to produce the second half of the expected string.
3.5.2 Effect of Task Snippet Position
Figure 4: Effect of key function position on one-step task with 5 distractors. Position is defined as index of the first token in the function normalized by the total number of tokens in the prompt, and grouped into 20 bins. 0 is the beginning of the prompt and 1 is at the end.
In the one-step retrieval task we see good performance with respect to key function position across the entire context window with two notable exceptions. At the 8k context size Mistral-7B-v0.1 and StarCoder2-7B are unable to perform the one-step task when the key function is more than ~4k tokens away from the generation site Figure 4. Both models use a sliding window attention mechanism with a 4k window size. While this mechanism theoretically allows the model to scale very large context sizes (Jiang et al. [4] reports a theoretical attention span of 131k tokens), our results indicate that the model fails to retrieve precise information at a distance greater than the sliding window.
Figure 5: Effect of key and value function position for two-step tasks. StarCoderBase-7B (top), StarCoder2-7B (bottom) with 5 distractors. Color represents accuracy@3 score.
The two-step task has two relevant task snippet positions, Figure 5 shows heatmaps of task performance vs. both key function position and value function position. We see that performance is worse when the key function appears before the value function in the prompt. We call this a forward reference and it turns out this has a large effect on performance. We also observe that performance is generally higher along the diagonal, i.e., when the functions are closer to each other.
For three-step and concatenation tasks, it becomes harder to directly analyze position effects, as there are three task-relevant code snippets with independent positions. We instead explore performance by forward references and spread between task-relevant snippets in the following sections.
3.5.3 Effect of Forward References
Figure 6: Effect of forward references by task variant, StarCoderBase-7B with 5 distractors. See Figure 9 in the appendix for all models and number of distractors.
We define a forward reference as a function calling a not-yet-defined function. We find that forward references have a strong negative impact on task performance. Figure 6 shows the performance of StarCoderBase-7B with respect to the number of forward references. Accuracy@3 score drops from 87% to 34% for two-step task at 4k context size when the order of the two task snippets is reversed. We find similar behavior with respect to forward references across all the models in our experiment (Figure 9 in Appendix A).
3.5.4 Effect of Task Snippet Spread
Figure 7: Effect of spread (normalized to the number of tokens in the context) for StarCoderBase-7B with 4k context size and 5 distractors. For data on other models and context sizes, see Figure 10, Figure 11, and Figure 12.
We saw in Figure 5 that the order of functions in multi-step tasks and the distance between functions affects performance. We define spread as the distance between the first token of the first task-related snippet and the last token of the last task-related snippet, not including the assert statement at the end of the prompt.
Figure 7 shows the effect of task snippet spread with StarCoderBase-7B with 4k context length. For the two-step task, as spread increases, the performance gap between zero and one forward reference grows, we find that this result is consistent across models Figure 10. For the three-step task a similar effect is clearly visible at 2k and 4k context size. However this effect is not consistent across different models Figure 11.
In the concatenation task, the effect is less pronounced as model’s performance is low regardless of spread and the number of forward references and there is not much room for separation Figure 12.
4 Improving Retrieval Performance with Call Graph Comments
We saw in Section 3.5.3 that performance degrades for the multi-step tasks in the presence of forward references. We considered if injecting information about function call relationships into the prompt would help performance. This information is easy to obtain from static analysis of code (or in our case at prompt construction time), and if beneficial, provides a lightweight approach to improving performance at inference time. We focus on our most difficult tasks, namely the three-step and concatenation retrieval task with 5 distractors using the same models as in our previous experiment.
4.1 Experiment Setup
We add comments above each function that contain a lightweight description of the call graph associated with that function. We can construct these comments by annotating which functions are “called by” other functions, or annotating which functions “call” a given function, or a combination of both. In the three-step retrieval task we can do this transitively and add annotation of all the functions that will be called to compute the result of a given function.
We considered two templates for these comments, in one variation (“names only”) we simply included a comma-separated list of function names. In the other (“full sentence”) we use the phrases: “This function is called by ” and “This function calls ” followed by the comma-separated list of function names. See Appendix G for examples.
4.2 Results
Table 2: Call graph comment performance across all models. Accuracy@3
(a) By template
Template Comment Types
Calls Called By Both
Full-Sentence 20.4 31.7 36.7
Names-Only 15.7 28.3 27.2
(b) By task (full sentence template only)
Task Comment Types
None Calls Called By Both
Three-step 32.1 39.1 64.6 68.5
Concatenation 4.5 9.2 11.9 17.7
We observe that the addition of call-graph comments has a positive effect on task performance. 2(b) shows a ~4x improvement on the concatenation task, and a ~2x improvement on the three-step retrieval task with the addition of call-graph comments aggregated across models. We do note that StarCoderBase-1B sees no improvement on the concatenation task. See Appendix B for detailed results.
Full sentence vs. function names only:
The full-sentence template performs better than the names-only version (2(a)). This suggests that the models are able to take some advantage of the more complete natural language description and that there may be room for further improvement by tuning the template. We focus the rest of our analysis on the full-sentence condition.
Call graph directions:
2(b) shows that while most of the benefit comes from the “X is called by Y” type of comment, using both directions produces the best improvement overall.
Table 3: Call graph comment performance by depth across all models. Accuracy@3
# Forward References Comment Type
None Next Hop Full
0 38.0 61.3 67.6
1 33.5 62.9 71.0
2 20.8 47.0 59.3
Full graph vs. next-hop comments:
We noted earlier that in the three-step case we add comments describing the full call graph (i.e., up to two steps away). We experimented with only adding comments about the next function in the call graph and found that the full graph version performs best (Table 3).
Figure 8: Effect of call graph comments on StarcoderBase-7B performance across context sizes, spread of task snippets and number of forward references.
Context size, spread, and forward references:
Figure 8 shows that call-graph comments improve performance across all context lengths, number of forward references, and task-snippet spreads for StarCoderBase-7B (this trend holds for other models where the call graph comments make a difference). Full plots for all the models are found in Appendix B. We note that with respect to spread for the concatenation task, most of the improvement appears when the task-relevant snippets are closer together.
5 Discussion
We found that models were able to perform well at the simpler one-step recall task over their entire context lengths (2(a)). However our experiments highlight the importance of distinguishing between long-context support and the ability to handle long-range dependencies. This is illustrated by models like Mistral and StarCoder2 that appear to make a trade-off between the two. While sliding window attention greatly expands their supported context sizes, our experiments show that when distance between relevant snippets is greater than the sliding window size used by these models, performance degrades quickly. Our tasks require precise recall over a long distance and it appears to be difficult to pass on information from earlier sliding windows in a way that allows successful completion of the task.
As different approaches to scaling attention to larger context windows proliferate, practitioners should take care to understand and mitigate these effects if their task relies on precise recall of long-range dependencies. In the context of code generation, builders of retrieval augmented generation systems may find it desirable to rewrite code to bring code snippets that depend on each other closer together.
Our multi-step tasks test the ability of models to reason through indirection introduced by simple yet common operations like function calling and concatenation. We find that models perform significantly worse on these tasks than the simpler one-step recall task (2(a)). This is particularly true in the presence of distractors, which we believe are important for appropriately pressure testing the ability of models to do the required reasoning (Figure 3).
We note that models we tested struggled even more in the presence of forward references (Figure 6). In this situation all the information in the prompt is the same, but is presented in a different order. Humans and compilers are able to resolve this problem and it would be ideal if models could do the same. While we do not know what causes this behaviour, one hypothesis is that the causal attention used during training favors attending backwards in the context window and propagating information from earlier to later tokens.
We found that adding simple natural-language text annotations that describe the call graph greatly improves performance on the more difficult multi-step tasks in all cases, including where forward references are present. Most of the benefit seems to come from just having the function name of the forward-reference appear first, though there also appears to be room to explore tuning the comment text to improve performance (2(b)).
These findings suggest a couple of avenues for practitioners to explore in constructing prompts for code generation. Firstly it may be possible to rewrite code to reduce forward references while maintaining the semantics of the code, and secondly that extra information that can be gleaned from existing methods such as static analysis can be injected as comments to help models reason over the code.
6 Limitations
Our experiments uses simple synthetically generated functions that are not as realistic as real code in the wild. Recommendations such as re-ordering functions to bring dependant code closer together or adding annotations like call-graph comments would need to be tested in real retrieval-augmented code generation environments. We do not do in-depth investigation of potential internal mechanisms that lead to the behaviours we observed, we describe some hypotheses but otherwise leave that to future work.
7 Acknowledgements
We thank Alan Leung, Titus Barik, Dominik Moritz, Samira Abnar, Forrest Huang, Jeff Nichols and Barry-John Theobald for their helpful feedback and support
Appendix A Detailed Results: Experiment 1
We include detailed tables of accuracy@3 scores (percent scale). Table 4 shows overall scores by task and context length (2k, 4k, and 8k); Table 5 shows scores by the number of distractors; and Table 6 shows scores by the number of forward references.
Table 4: Overall accuracy@3 scores for each model, task, and context size. Each score in the context length columns is the mean across all three distractor conditions. The “mean” column shows the average score for the three context sizes.
Task Model 2k 4k 8k mean
one-step Mistral-7B-v0.1 99.8±0.1 99.8±0.1 46.4±0.9 61.5
StarCoder2-7B 100.0±0.0 99.8±0.1 68.1±0.8 67.0
StarCoderBase-1B 97.1±0.3 98.3±0.2 98.5±0.1 73.5
StarCoderBase-7B 100.0±0.0 99.9±0.0 99.6±0.1 74.9
StarCoderBase-15.5B 99.8±0.1 99.5±0.1 99.0±0.1 74.6
two-step Mistral-7B-v0.1 63.1±0.7 58.3±0.6 16.7±0.5 34.6
StarCoder2-7B 81.9±0.5 82.6±0.5 46.4±0.6 52.7
StarCoderBase-1B 73.3±0.6 73.5±0.5 73.7±0.5 55.1
StarCoderBase-7B 81.6±0.5 80.1±0.5 79.2±0.5 60.2
StarCoderBase-15.5B 82.2±0.5 74.7±0.5 64.0±0.6 55.2
three-step Mistral-7B-v0.1 53.1±0.6 41.5±0.5 14.7±0.4 27.3
StarCoder2-7B 78.9±0.5 75.3±0.5 40.0±0.6 48.6
StarCoderBase-1B 55.9±0.6 55.2±0.6 52.9±0.6 41.0
StarCoderBase-7B 67.0±0.6 66.9±0.5 66.2±0.6 50.0
StarCoderBase-15.5B 65.8±0.6 57.1±0.6 45.3±0.6 42.1
concatenation Mistral-7B-v0.1 17.3±0.4 6.7±0.3 0.7±0.1 6.2
StarCoder2-7B 34.7±0.6 18.3±0.5 4.5±0.3 14.4
StarCoderBase-1B 1.7±0.1 0.6±0.1 0.1±0.0 0.6
StarCoderBase-7B 21.9±0.5 13.4±0.4 8.0±0.3 10.8
StarCoderBase-15.5B 34.4±0.6 21.6±0.5 15.2±0.4 17.8
Table 5: Overall model performance vs. number of distractors.
One-step Task
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 100.0 99.9 99.6 100.0 100.0 99.5 48.5 48.2 42.8 StarCoder2-7B 100.0 100.0 99.9 100.0 100.0 99.6 76.3 69.3 59.8 StarCoderBase-1B 100.0 99.1 93.2 99.9 98.4 96.7 99.9 98.8 97.0 StarCoderBase-7B 100.0 100.0 99.9 100.0 100.0 99.9 100.0 99.7 99.1 StarCoderBase-15.5B 99.9 99.8 99.8 100.0 99.5 99.0 99.8 99.3 98.1
Two-step Task
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 87.4 74.5 40.6 76.7 63.9 34.4 22.7 18.9 8.6 StarCoder2-7B 99.2 85.5 65.5 97.9 86.8 63.0 62.6 49.1 27.5 StarCoderBase-1B 98.5 80.9 47.3 96.2 82.5 41.8 89.5 79.2 52.4 StarCoderBase-7B 99.2 86.3 64.0 97.9 81.9 60.6 95.4 85.3 56.8 StarCoderBase-15.5B 93.6 80.9 74.4 89.6 68.8 65.8 82.0 64.3 45.8
Three-step Task
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 81.4 54.0 23.8 61.4 42.3 21.0 19.5 14.7 10.0 StarCoder2-7B 95.6 83.0 58.2 94.5 85.4 46.0 60.0 39.5 20.5 StarCoderBase-1B 80.7 59.0 28.1 70.9 68.3 26.4 79.6 58.7 20.5 StarCoderBase-7B 91.7 68.0 41.2 82.6 75.9 42.3 88.0 77.1 33.4 StarCoderBase-15.5B 81.4 66.5 49.5 69.7 62.1 39.6 65.0 49.5 21.5
Concatenation Task
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 27.0 18.5 6.4 12.8 6.2 1.2 1.1 0.6 0.4 StarCoder2-7B 54.3 36.2 13.4 26.5 23.0 5.4 8.6 3.9 1.1 StarCoderBase-1B 2.9 1.5 0.6 1.1 0.5 0.2 0.3 0.1 0.0 StarCoderBase-7B 37.5 19.2 9.1 26.9 11.7 1.6 16.1 7.3 0.5 StarCoderBase-15.5B 50.5 36.6 16.1 35.6 21.9 7.3 27.5 13.3 4.6
Table 6: Overall model performance vs. number of forward references. 5 distractors. Note that forward references cannot exist in one-step tasks.
Two-step Task
Context Size 2k 4k 8k # Forward References 0 1 0 1 0 1 Mistral-7B-v0.1 60.1 21.1 46.4 22.5 8.0 9.2 StarCoder2-7B 73.5 57.6 69.4 56.5 28.1 26.8 StarCoderBase-1B 55.3 39.4 49.7 33.9 65.6 39.3 StarCoderBase-7B 90.0 38.1 87.2 34.1 76.0 37.7 StarCoderBase-15.5B 85.5 63.4 82.3 49.2 63.4 28.3
Three-step Task
Context Size 2k 4k 8k # Forward References 0 1 2 0 1 2 0 1 2 Mistral-7B-v0.1 28.1 24.4 16.8 19.8 22.3 16.9 2.7 11.7 10.3 StarCoder2-7B 61.5 60.5 45.7 46.2 48.1 37.5 14.4 23.1 16.0 StarCoderBase-1B 32.3 29.5 18.1 29.5 28.5 14.9 19.8 23.8 8.2 StarCoderBase-7B 63.4 41.3 18.8 65.4 42.0 20.5 41.9 35.6 15.9 StarCoderBase-15.5B 57.2 50.6 37.2 54.0 39.7 24.9 34.5 21.0 10.5
Concatenation Task
Context Size 2k 4k 8k # Forward References 0 1 2 0 1 2 0 1 2 Mistral-7B-v0.1 11.8 4.3 3.1 2.4 0.6 0.6 0.8 0.2 0.2 StarCoder2-7B 17.7 14.7 7.9 6.4 5.4 4.2 1.8 0.7 0.7 StarCoderBase-1B 1.6 0.3 0.0 0.5 0.1 0.0 0.0 0.0 0.0 StarCoderBase-7B 13.9 9.3 4.0 2.0 1.6 1.2 0.6 0.6 0.3 StarCoderBase-15.5B 22.9 14.2 11.0 10.1 6.1 5.8 6.8 4.5 2.6
Figure 9: Effect of forward references by task variant, with 5 distractors. The data is showin Table 6
Figure 10: Effect of spread for two step tasks for each model, with 5 distractors.
Figure 11: Effect of spread for three step tasks for each model, with 5 distractors.
Figure 12: Effect of spread for concatenation tasks for each model, with 5 distractors.
Mistral-7B-v0.1
StarCoder2-7B
StarCoderBase-1B
StarCoderBase-7B
StarCoderBase-15.5B
Figure 13: Effect of key and value function position for two-step tasks, with 5 distractors. Color represents accuracy@3 score.
Appendix B Detailed Results: Call Graph Comment Experiment
We include details of model performance (accuracy@3 percent scores) for the full-sentence template variant of the call graph experiment.
Table 7: Call graph comment performance by model.
Task Model Comment Type
None Calls Called-By Both
Three Step Mistral-7B-v0.1 18.2 16.7 33.8 39.6
StarCoder2-7B 41.6 57.4 80.2 90.8
StarCoderBase-1B 25.0 30.5 46.5 47.4
StarCoderBase-7B 39.0 48.7 86.0 87.2
StarCoderBase-15.5B 36.9 49.5 85.3 87.5
Concatenation Mistral-7B-v0.1 2.7 3.8 5.6 8.9
StarCoder2-7B 6.6 14.3 27.4 31.2
StarCoderBase-1B 0.3 0.3 0.3 0.4
StarCoderBase-7B 3.7 17.6 16.0 30.2
StarCoderBase-15.5B 9.3 20.0 33.7 40.6
Table 8: Next-hop vs. full graph comments performance on three-step retrieval.
Model # Forward References Comment Type
None Next Hop Full Graph
Mistral-7B-v0.1 0 16.8 39.6 53.4
1 19.5 29.9 39.8
2 14.7 17.1 25.3
StarCoder2-7B 0 40.7 83.2 90.9
1 43.9 80.5 92.5
2 33.1 61.1 84.3
StarCoderBase-1B 0 27.2 28.4 32.9
1 27.3 46.3 52.1
2 13.8 29.9 43.4
StarCoderBase-7B 0 56.9 88.7 92.5
1 39.6 78.9 91.2
2 18.4 46.8 66.0
StarCoderBase-15.5B 0 48.6 66.6 76.6
1 37.1 78.7 89.3
2 24.2 80.2 91.5
Figure 14: Effect of call graph comments on Mistral-7B-v0.1.
Figure 15: Effect of call graph comments on Starcoder2-7B performance.
Figure 16: Effect of call graph comments on StarCoderBase-1B performance.
Figure 17: Effect of call graph comments on StarCoderBase-7B performance.
Figure 18: Effect of call graph comments on StarCoderBase-15.5B performance.
Appendix C Hyperparameters for Generation
For all experiments we use nucleus sampling with the following parameters.
Table 9: Generation hyperparameters.
Hyperparameter Value
Temperature 0.8
Top p 0.95
Top k 0
Batch size 1
Output samples per input prompt 10
Appendix D Confidence Intervals
All confidence intervals (error bars and plus/minus values) are calculated as 95% confidence intervals of the estimated mean following assumptions of normal distribution and independent samples.
Appendix E Compute Information
We ran all experiments on machines with a single A100 GPU with 80GB of VRAM on a cloud provider. Experiments were run over the course of about a month and we estimate ~1000 GPU hours were used including runs that had bugs in them and additional experiments that are not part of the paper. We use FlashAttention [6] to improve memory usage and latency of generation.
Appendix F Prompt Samples for Multi Step Key Retrieval
def iaizjb_184360_440195():
return "wxmbrnpokw"
def string_to_md5(text):
Given a string ’text’, return its md5 hash equivalent string.
If ’text’ is an empty string, return None.
$>$ string_to_md5(’Hello world’) == ’3e25960a79dbc69b674cd4ec67a72c62’
import hashlib
return hashlib.md5(text.encode(’ascii’)).hexdigest() if text else None
def egllun_467846():
return "eooyfwmxln"
def count_upper(s):
Given a string s, count the number of uppercase vowels in even indices.
For example:
count_upper(’aBCdEf’) returns 1
count_upper(’abcdefg’) returns 0
count_upper(’dBBE’) returns 0
count = 0
for i in range(0,len(s),2):
if s[i] in "AEIOU":
count += 1
return count
def solution(lst):
"""Given a non-empty list of integers, return the sum of all of the odd elements that are in even positions.
Examples
solution([5, 8, 7, 1]) ==> 12
solution([3, 3, 3, 3, 3]) ==> 9
solution([30, 13, 24, 321]) ==>0
return sum([x for idx, x in enumerate(lst) if idx%2==0 and x%2==1])
def choose_num(x, y):
"""This function takes two positive numbers x and y and returns the
biggest even integer number that is in the range [x, y] inclusive. If
there’s no such number, then the function should return -1.
For example:
choose_num(12, 15) = 14
choose_num(13, 12) = -1
if x > y:
return -1
if y % 2 == 0:
return y
if x == y:
return -1
return y - 1
def digitSum(s):
"""Task
Write a function that takes a string as input and returns the sum of the upper characters only’
ASCII codes.
Examples:
digitSum("") => 0
digitSum("abAB") => 131
digitSum("abcCd") => 67
digitSum("helloE") => 69
digitSum("woArBld") => 131
digitSum("aAaaaXa") => 153
if s == "": return 0
return sum(ord(char) if char.isupper() else 0 for char in s)
def multiply(a, b):
"""Complete the function that takes two integers and returns
the product of their unit digits.
Assume the input is always valid.
Examples:
multiply(148, 412) should return 16.
multiply(19, 28) should return 72.
multiply(2020, 1851) should return 0.
multiply(14,-15) should return 20.
return abs(a % 10) * abs(b % 10)
def right_angle_triangle(a, b, c):
Given the lengths of the three sides of a triangle. Return True if the three
sides form a right-angled triangle, False otherwise.
A right-angled triangle is a triangle in which one angle is right angle or
90 degree.
Example:
right_angle_triangle(3, 4, 5) == True
right_angle_triangle(1, 2, 3) == False
return a*a == b*b + c*c or b*b == a*a + c*c or c*c == a*a + b*b
def add_elements(arr, k):
Given a non-empty array of integers arr and an integer k, return
the sum of the elements with at most two digits from the first k elements of arr.
Example:
Input: arr = [111,21,3,4000,5,6,7,8,9], k = 4
Output: 24 # sum of 21 + 3
Constraints:
1. 1 <= len(arr) <= 100
2. 1 <= k <= len(arr)
return sum(elem for elem in arr[:k] if len(str(elem)) <= 2)
def vskfby_510934():
return "thwtyqwjws"
def qgtsin_336194_iwdghb():
return iaizjb_184360_440195()
def awdpgq_293061_vwetvu():
return rbwofb_803321_331141()
def even_odd_count(num):
"""Given an integer. return a tuple that has the number of even and odd digits respectively.
Example:
even_odd_count(-12) ==> (1, 1)
even_odd_count(123) ==> (1, 2)
even_count = 0
odd_count = 0
for i in str(abs(num)):
if int(i)%2==0:
even_count +=1
else:
odd_count +=1
return (even_count, odd_count)
def minSubArraySum(nums):
Given an array of integers nums, find the minimum sum of any non-empty sub-array
of nums.
Example
minSubArraySum([2, 3, 4, 1, 2, 4]) == 1
minSubArraySum([-1, -2, -3]) == -6
max_sum = 0
s = 0
for num in nums:
s += -num
if (s < 0):
s = 0
max_sum = max(s, max_sum)
if max_sum == 0:
max_sum = max(-i for i in nums)
min_sum = -max_sum
return min_sum
def qvkrlp_701464_787942():
return vskfby_510934()
def rdcxoi_135343():
return egllun_467846()
def iscube(a):
Write a function that takes an integer a and returns True
if this ingeger is a cube of some integer number.
Note: you may assume the input is always valid.
Examples:
iscube(1) ==> True
iscube(2) ==> False
iscube(-1) ==> True
iscube(64) ==> True
iscube(0) ==> True
iscube(180) ==> False
a = abs(a)
return int(round(a ** (1. / 3))) ** 3 == a
def special_factorial(n):
"""The Brazilian factorial is defined as:
brazilian_factorial(n) = n! * (n-1)! * (n-2)! * ... * 1!
where n > 0
For example:
$>$ special_factorial(4)
288
The function will receive an integer as input and should return the special
factorial of this integer.
fact_i = 1
special_fact = 1
for i in range(1, n+1):
fact_i *= i
special_fact *= fact_i
return special_fact
def is_equal_to_sum_even(n):
"""Evaluate whether the given number n can be written as the sum of exactly 4 positive even numbers
Example
is_equal_to_sum_even(4) == False
is_equal_to_sum_even(6) == False
is_equal_to_sum_even(8) == True
return n%2 == 0 and n >= 8
assert rdcxoi_135343() ==
Listing 1: Sample prompt for two-step retrieval task with 2000 token context size and 5 distractors.
Appendix G Prompt Samples for Call Graph Comments Experiment
# This function calls oswgtr_325169_862229
# This function is called by aokwfl_208971_hwmofh
def gqpvbp_138573():
return oswgtr_325169_862229()
# This function calls gqpvbp_138573 and oswgtr_325169_862229
def aokwfl_208971_hwmofh():
return gqpvbp_138573()
# This function calls lhezee_508969 and hjdnwl_724283
def oftoyy_286138():
return lhezee_508969()
# This function is called by gqpvbp_138573 and aokwfl_208971_hwmofh
def oswgtr_325169_862229():
return "kyfgholcrg"
# This function calls gjobme_651008_tymmij
# This function is called by wwzfoa_904885
def bweckw_860527_nykiyp():
return gjobme_651008_tymmij()
# This function is called by lhezee_508969 and oftoyy_286138
def hjdnwl_724283():
return "pwincnzyqh"
# This function is called by bweckw_860527_nykiyp and wwzfoa_904885
def gjobme_651008_tymmij():
return "axxtrhucug"
# This function calls bweckw_860527_nykiyp and gjobme_651008_tymmij
def wwzfoa_904885():
return bweckw_860527_nykiyp()
Listing 2: Full sentence call graph comments for three step retrieval. 5 distractors. HumanEval functions excluded for brevity.
# oswgtr_325169_862229
# aokwfl_208971_hwmofh
def gqpvbp_138573():
return oswgtr_325169_862229()
# gqpvbp_138573, oswgtr_325169_862229
def aokwfl_208971_hwmofh():
return gqpvbp_138573()
# lhezee_508969, hjdnwl_724283
def oftoyy_286138():
return lhezee_508969()
# gqpvbp_138573, aokwfl_208971_hwmofh
def oswgtr_325169_862229():
return "kyfgholcrg"
# gjobme_651008_tymmij
# wwzfoa_904885
def bweckw_860527_nykiyp():
return gjobme_651008_tymmij()
# lhezee_508969, oftoyy_286138
def hjdnwl_724283():
return "pwincnzyqh"
# bweckw_860527_nykiyp, wwzfoa_904885
def gjobme_651008_tymmij():
return "axxtrhucug"
# bweckw_860527_nykiyp, gjobme_651008_tymmij
def wwzfoa_904885():
return bweckw_860527_nykiyp()
Listing 3: Function names only call graph comments for three step retrieval. 5 distractors. HumanEval functions excluded for brevity.